OpenStudy (anonymous):

the function given by P(d) =1+d/33 gives the pressure, in atmosopheres, at a depth of d feet. for what depths id the pressure atleast 1 atmosphere and at most 7

jimthompson5910 (jim_thompson5910):

1 <= P(d) <= 7 1 <= 1 + d/33 <= 7 1 - 1 <= d/33 <= 7 - 1 0 <= d/33 <= 6 0*33 <= d <= 6*33 0 <= d <= 198 So for depths between 0 ft and 198 ft, the pressure will be at least 1 atmosphere and at most 7 atmospheres.

OpenStudy (anonymous):

jim my friend, you are the man